OpenStudy (anonymous):

Let z_1=2+3i and z_2=1-i. Perform the indicated operations and write the solutions in the form a+bi. |z_2|= ?

OpenStudy (anonymous):

a. -1-i b. 1+i c. -1+i d. sqrt2

OpenStudy (anonymous):

Underscore refers to base

OpenStudy (amorfide):

what do you mean by base?

OpenStudy (anonymous):

z_1= z and then a little one right beside the z

OpenStudy (amorfide):

okay so  this means the modulus of z_2 modulus is Pythagoras so you do 1² + (-i)²=1 + i²=1-1 1-1=0 root 0=0 modulus=0 i assume this is what you mean? unless the base means something different lol sorry

OpenStudy (anonymous):

Yes this is what I mean. Thanks! But I don't think it matches with any of the options :\

OpenStudy (anonymous):

modulus is sqrt((1^2)+(-1)^2)

OpenStudy (anonymous):

simplify that...

OpenStudy (anonymous):

Simplified to sqrt2! Thanks!

OpenStudy (amorfide):

my bad i squared i not the imaginary part of i sorry

OpenStudy (anonymous):

I wish I could give more than one medal. Thanks amorfide for the help and explanation!

OpenStudy (anonymous):

we'll split it... :)